OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? y = 4x − 1 y = 4x − 5 y = 4x − 10 y = 4x + 1

OpenStudy (anonymous):

y= 4x - 10

OpenStudy (anonymous):

Thanks! Mind helping me on a bit more?

OpenStudy (anonymous):

sure. .

OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 3) and (2, 7)? x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

OpenStudy (anonymous):

x - y =-5

OpenStudy (anonymous):

Thank you! Here's the last one: Consider the functions f(x) = 2x2 + 1 and g(x) = x + 2. What is the value of f[g(3)]? 101 21 11 51

OpenStudy (anonymous):

51

OpenStudy (anonymous):

You're a life-saver. :)

OpenStudy (anonymous):

glad to help. .